What current, in units of milliAmperes, would produce a potential difference of 7.5 V across a 25 Ω resistor?
1
Expert's answer
2018-03-09T10:59:07-0500
From ohms law, the current flowing through a conductor is directly proportional to the voltage across the conductor.
V=IR
therefore, I=V/R. =7.5V/25Ω=0.3A
0.3×1000=300mA